OpenStudy (anonymous):

The first term of a sequence is 1.If each term after the first is the product of -2 and the preceding term,what is the sixth term of the sequence?

OpenStudy (gorv):

1 first -2*1=-2 second -2*(-2)=4 third 4*(-2)=-8 fourth

OpenStudy (gorv):

similarly found other